OpenStudy (anonymous):

is the following an arithmetic sentence 5,0,-5,-10 plz help

OpenStudy (ajprincess):

u mean sequence?

OpenStudy (ajprincess):

@Kayleekay

OpenStudy (anonymous):

0=5+(-5) -5=0+(-5) -10=-5 +(-5) What can you conclude?